OpenStudy (anonymous):

Can someone please explain how to solve 1-n/5 = 12.

OpenStudy (midhun.madhu1987):

 is the question like this?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ill send it to you

OpenStudy (anonymous):

Thank you

OpenStudy (midhun.madhu1987):

Just subtract 1 from both sides.. what u get?

OpenStudy (anonymous):

I got n/5=11

OpenStudy (midhun.madhu1987):

-n/5 = 11 right?

OpenStudy (midhun.madhu1987):

Now multiply both sides by 5.. wht u get?

OpenStudy (anonymous):

Step 1: Simplify both sides of the equation -1/5n + 1 = 12 Step 2: Subtract 1 from both sides -1/5n =11 Step 3: Divide both sides by (-1)/5.....-1/5n/-1/5= 11/-1/5 Answer n= -55

OpenStudy (anonymous):

Does this help?

OpenStudy (anonymous):

Yes thanks